OpenStudy (anonymous):

Solve the proportion. -5/7=15/x

OpenStudy (astrophysics):

\[\frac{ -5 }{ 7 }=\frac{ 15 }{ x }\] Cross multiply

OpenStudy (astrophysics):

\[-5x=15(7)\]

OpenStudy (astrophysics):

\[-5x=105 \implies x = \frac{ -105 }{ 5 } \implies x = -21\]

OpenStudy (anonymous):

Thank you!!!